OpenStudy (anonymous):

3 8 ---- + ---- add and simplify yv^2 y^2v

OpenStudy (anonymous):

\[\frac{3}{(y) v^2}+\frac{8}{\left(y^2\right) v} \]Is this the problem statement?

OpenStudy (anonymous):

it is how i wrote it

OpenStudy (anonymous):

Cross multiply and simplify\[\frac{3}{(y) v^2}+\frac{8}{\left(y^2\right) v}=\frac{3v y^2+8v^2 y}{v y^2 v^2 y}=\frac{8 v+3 y}{v^2 y^2} \]